OpenStudy (mariamikayla):

f(f(x)) = x^2 -x +1 , what is f(x) ?

OpenStudy (rolldextro):

They didnt define f(x) for you?

OpenStudy (mariamikayla):

f(x) i guess it's a 2nd degree function

OpenStudy (rolldextro):

Are you sure its not f(x) instead of f(f(x))?

OpenStudy (mariamikayla):

yes

OpenStudy (mariamikayla):

it's f(f(x)) which means fof(x)

OpenStudy (rolldextro):

I know I just wanted to make sure

OpenStudy (rolldextro):

Assuming this is what you mean. f(x) = x^2-x+1 then f(f(x))=x^4 - 2 x^3 + 2 x^2 - x + 1

OpenStudy (rolldextro):

Just take your function and replace every x with it and simplify.

OpenStudy (mariamikayla):

it's inverse...

OpenStudy (mariamikayla):

you have to find the function,it's not given

OpenStudy (rolldextro):

(1 ± Sqrt[-3 + 4 x])/2

OpenStudy (mariamikayla):

what is that?

OpenStudy (rolldextro):

You will get two functions. one will be (1 + Sqrt[-3 + 4 x])/2 and the other one will be (1 - Sqrt[-3 + 4 x])/2 depends on what you need